### v1

# Gromov-Hausdorff convergence of metric spaces of UCP maps

Abstract.
It is shown that van Suijlekom’s technique of imposing a set of conditions on operator system spectral triples ensures Gromov-Hausdorff convergence of sequences of sets of unital completely positive maps (equipped with the BW-topology which is metrizable). This implies that even when only a part of the spectrum of the Dirac operator is available together with a certain truncation of the $C^{*}$-algebra, information about the geometry can be extracted.
2020 Mathematics Subject Classification: 46L07, 46L87, 47C15, 47L25, 58B34.
Keywords: Operator system, UCP maps, Spectral truncation, Gromov–Hausdorff convergence
Acknowledgement: The authors are thankful to Walter van Suijlekom whose India visit under the Jubilee Chair Professorship of the Indian of Academy of Sciences was pivotal for this paper.
1. Introduction
Let $C(S^{1})$ be the commutative $C^{*}$-algebra of all continuous functions on the unit circle $S^{1}$. The Fejer-Riesz operator system $C(S^{1})_{(n)}$ for any $n\geq 1$ is the vector space of all continuous functions $f$ on the unit circle $S^{1}$ such that the Fourier coefficients $\hat{f}(k)$ vanish for all $k\in\mathbb{Z}$ with $|k|\geq n$, i.e., $f\in C(S^{1})_{(n)}$ if and only if
$f(z)=\sum_{k=-(n-1)}^{n-1}a_{k}z^{k}$
for $z\in S^{1}$. As a subspace of the $C^{*}$-algebra $C(S^{1})$, it is *-closed because the involution is given by
$\sum_{k=-(n-1)}^{n-1}a_{k}z^{k}\rightarrow\sum_{k=-(n-1)}^{n-1}\overline{a_{-k}}z^{k}$
and contains the identity of the $C^{*}$-algebra. So, it is an operator system.
On the other hand, the $n\times n$ Toeplitz matrices over $\mathbb{C}$, denoted by $\mathcal{T}_{n}$ form an operator system because it is a *-closed unital subspace of the $C^{*}$-algebra of all $n\times n$ matrices over $\mathbb{C}$.
Consider the linear map $\delta:\mathcal{T}_{n}\rightarrow C(S^{1})_{(n)}$ which takes a Toeplitz matrix $t=((\tau_{k-l}))_{k,l=0}^{n-1}$ to the linear functional $\varphi_{t}$ on $C(S^{1})_{(n)}$ defined by
$\varphi_{t}(f)=\sum_{k=-(n-1)}^{n-1}\tau_{-k}a_{k}$
where $f(z)=\sum_{k=-(n-1)}^{n-1}a_{k}z^{k}$. Connes and van Suijlekom in Proposition 4.6 of [2] showed that this is a linear unital order isomorphism between the dual of $C(S^{1})_{(n)}$ and $\mathcal{T}_{n}$. In Theorem 1.1 in [3], Farenick showed that this duality identification is unital completely isometrically isomorphic.
We are motivated by the convergence results in the literature, e.g., the Gromov-Hausdorff convergence of quantum metric spaces in the seminal paper [10]. More recently, in [2] and [13], state spaces of the operator systems $C(S^{1})_{(n)}$ and $\mathcal{T}_{n}$ were dealt with. It was shown that in spite of the duality mentioned above, the sequences of state spaces $\mathcal{S}(C(S^{1})_{(n)})$ and $\mathcal{S}(\mathcal{T}_{n})$ both converge to the state space $\mathcal{S}(C(S^{1}))$. More generally, in [13], general conditions were found on operator system spectral triples that guarantee the Gromov-Hausdorff convergence of state spaces.
Given an operator system $E$ and a (complex, separable) Hilbert space $\mathcal{K}$, let $\mathcal{UCP}_{\mathcal{K}}(E)$ denote the collection of all unital completely positive (UCP) maps from $E$ into the algebra of all bounded operators $\mathcal{B}(\mathcal{K})$. We shall show in this note that the general conditions due to van Suijlekom mentioned above actually guarantee much more, viz., the Gromov-Hausdorff convergence of sequences of sets of UCP maps. In particular, $\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ and $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ both converge in the Gromov-Hausdorff metric to $\mathcal{UCP}_{\mathcal{K}}(C(S^{1}))$ as $n$ tends to infinity, implying that for large $n$, $\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ and $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ are close in the Gromov-Hausdorff metric.
The main result is Theorem 1 where we consider a sequence of operator system spectral triples $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ related to an operator system spectral triple $(\mathcal{E},\mathcal{H},D)$ by a $\mathrm{C}^{1}$-approximate complete order isomorphism. Section 4 contains applications of the theorem to examples like the ones mentioned above. The $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ are often called spectral truncations of $(\mathcal{E},\mathcal{H},D)$. Spectral truncations have a widespread appeal - see [4], [8], [13] and [11] for noncommutative geometry, [7] for compact quantum groups, and [5] for machine learning.
2. Sets of CP maps as metric spaces
2.1. The BW-topology
To talk of the Gromov-Hausdorff metric, we need to make a metric space out of $\mathcal{UCP}_{\mathcal{K}}(E)$ first. For the following, a good reference is [9].
Definition 2.1 (BW-topology).
For Banach spaces $X$ and $Y$, let $\mathcal{B}(X,Y^{*})$ denote the bounded linear maps of $X$ into $Y^{*}$. For fixed vectors $x$ in $X$ and $y$ in $Y$, define a linear functional $x\otimes y$ on $\mathcal{B}(X,Y^{*})$ by $x\otimes y(L)$ = $L(x)(y)$. Let $Z$ denote the closed linear span in $\mathcal{B}(X,Y^{*})$ of these elementary tensors. Then $\mathcal{B}(X,Y^{*})$ is isometrically isomorphic to $Z^{*}$ with the duality given by
$\langle L,x\otimes y\rangle=x\otimes y(L)$
The weak topology that is induced on $\mathcal{B}(X,Y^{*})$ by this identification is called the BW topology ((((for bounded weak)))).
If $Y$ is the algebra $\mathcal{B}(\mathcal{H})$ of bounded operators on a Hilbert space $\mathcal{H}$, then a bounded net $L_{\lambda}$ in $\mathcal{B}(X,\mathcal{B}(H))$ converges in the BW topology to $L$ if and only if $\langle L_{\lambda}(x)h,h^{\prime}\rangle$ converges to $\langle L(x)h,h^{\prime}\rangle$ for all $h$, $h^{\prime}$ in $\mathcal{H}$ and $x$ in $X$. Naturally, $\mathcal{UCP}_{\mathcal{K}}(E)$ inherits the BW-topology from $\mathcal{B}(E,\mathcal{B}(K))$. It is known that $\mathcal{UCP}_{\mathcal{K}}(E)$ is compact in the BW-topology [9, Theorem 7.4].
2.2. A metric for the BW-topology when $E=C(\Omega)$.
Let $(\Omega,d_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega})$ a compact metric space. Consider the set of all Lipschitz functions
$\text{Lip}(\Omega):=\{f:\Omega\to\mathbb{C}\text{ such that }\|f\|_{1}:=\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{d_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega}(x,y)}<\infty\}.$ (2.1)
Let $\{k_{n}\in\mathcal{K}:n\in\mathbb{N}\}$ be a fixed dense subset of the closed unit ball of $\mathcal{K}$. Define $d:\mathcal{UCP}_{\mathcal{K}}(C(\Omega))\times\mathcal{UCP}_{\mathcal{K}}(C(\Omega))\to[0,\infty)$ by
$d(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\text{Lip}(\Omega),\|f\|\leq 1,\|f\|_{1}\leq 1\right\}.$ (2.2)
Lemma 2.2.
The function $d$ given by (2.2) forms a metric on $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$.
Proof.
Clearly, symmetry and that the triangle inequality hold. Also, the implication $\varphi=\psi\implies d(\varphi,\psi)=0$ is immediate. Suppose now $d(\varphi,\psi)=0$. Then for all $f\in\text{Lip}(\Omega)$ and for all $p$,
$\left\langle(\varphi-\psi)(f)k_{p},k_{p}\right\rangle=0\implies\left\langle(\varphi-\psi)(f)k,k\right\rangle=0$ (2.3)
for all $k\in\mathcal{K}$. For $h,k\in\mathcal{K}$, (2.3) and the polarization identity
$\left\langle(\varphi-\psi)(f)h,k\right\rangle=\sum_{l=0}^{3}\frac{i^{l}}{4}\left\langle(\varphi-\psi)(f)(h+i^{l}k),h+i^{l}k\right\rangle$
together yield $\left\langle(\varphi-\psi)(f)h,k\right\rangle=0$. Hence $(\varphi-\psi)(f)=0$ for all $f\in\text{Lip}(\Omega)$, and consequently $\varphi=\psi$. This completes the proof.
Lemma 2.3.
The BW-topology on $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$ coincides with the metric topology on $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$ induced by the metric $d$ defined in (2.2).
Proof.
Let $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),d)$ and $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$ denote the metric topology and the BW topology on $\mathcal{UCP}_{K}\big{(}E\big{)}$, respectively. To show that these topologies are the same, it is enough to show that the identity map $I:(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)\to(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),d)$ is a homeomorphism. However, since $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$ is a compact space and $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),d)$ is Hausdorff, it suffices to show the continuity of $I$.
Let $\{\varphi_{\lambda}\}_{\lambda\in\Lambda}$ be a net in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$ such that $\varphi_{\lambda}$ converges to some $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$. Our claim is that $\varphi_{\lambda}$ also converges to $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),d)$. Suppose not. Then there exist $\varepsilon>0$ and an infinite subset $\hat{\Lambda}\subseteq\Lambda$ such that $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for all $\lambda\in\hat{\Lambda}$. Without loss of generality, we may assume that $\hat{\Lambda}=\mathbb{N}$. Thus, we have $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for $\lambda\in\mathbb{N}$. Therefore, there exists a sequence $\{f_{\lambda}\}$ in $\text{Lip}(X)$ with $\|f_{\lambda}\|\leq 1$ and $\|f_{\lambda}\|_{1}\leq 1$ such that
$\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|>2\varepsilon.$ (2.4)
Since $\|f_{\lambda}\|_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}\leq 1$, so $\{f_{\lambda}\}$ is a sequence of equicontinuous functions. Next, we also note that the series in (2.4) is convergent uniformly in $\lambda$ as
$\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|\leq 2.$
Hence, for the given $\varepsilon>0$, there exist a $N_{0}\in\mathbb{N}$ such that
$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p-1}}<\varepsilon.$
Consequently, for every $\lambda$
$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|<\varepsilon.$ (2.5)
Therefore, (2.4) and (2.5) together give
$\sum_{p=1}^{N_{0}}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|>\varepsilon.$ (2.6)
Since there are only finitely many $p$ in (2.6), we can choose a $p_{0}\in\{1,\ldots,N_{0}\}$ such that
$\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p_{0}},k_{p_{0}}\right\rangle\right|>\frac{\varepsilon}{N_{0}}.$ (2.7)
Since $\{f_{\lambda}\}$ is a uniformly bounded sequence in $E$ of equicontinuous functions, using the Arzelà-Ascoli theorem there exists a convergent subsequence (which we still denote by $\{f_{\lambda}\}$) that converges to some $f\in E$. Therefore, there exists an $M\in\mathbb{N}$ such that
$\|f_{\lambda}-f\|<\frac{\varepsilon}{4N_{0}}\text{ for all }\lambda\geq M.$ (2.8)
Finally, by (2.6), (2.7), and (2.8), we have
$\left|\left\langle(\varphi_{\lambda}-\varphi)(f)k_{p_{0}},k_{p_{0}}\right\rangle\right|$
$=$ $\left|\left(\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p_{0}},k_{p_{0}}\right\rangle\right)+\left(\left\langle\varphi(f_{\lambda}-f)k_{p_{0}},k_{p_{0}}\right\rangle+\left\langle\varphi_{\lambda}(f-f_{\lambda})k_{p_{0}},k_{p_{0}}\right\rangle\right)\right|$
$\geq$ $\Big{|}\left|\left(\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p_{0}},k_{p_{0}}\right\rangle\right)\right|-\left|\left(\left\langle\varphi(f_{\lambda}-f)k_{p_{0}},k_{p_{0}}\right\rangle+\left\langle\varphi_{\lambda}(f-f_{\lambda})k_{p_{0}},k_{p_{0}}\right\rangle\right)\right|\Big{|}$
$>$ $\frac{\varepsilon}{N_{0}}-\frac{2\varepsilon}{4N_{0}}=\frac{\varepsilon}{2N_{0}},$
which contradicts the fact that $\{\varphi_{\lambda}\}$ converges to $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$. Hence, $I$ is continuous. This completes the proof. ∎
2.3. A metric for the BW-topology when $E$ is a finite-dimensional operator system
The metric in (2.2) depended on the class in (2.1). In an operator system, we do not have access to such a class in general. To compensate for that loss, we need a Dirac operator $D$.
Definition 2.4.
An operator system spectral triple is a triple $(\mathcal{E},\mathcal{H},D)$ where $\mathcal{E}$ is a dense subspace of an operator system $E$ in $\mathcal{B}(\mathcal{H})$, $\mathcal{H}$ is a Hilbert space, and $D$ is a self-adjoint operator in $\mathcal{H}$ with compact resolvent, such that $[D,T]$ is a bounded operator for all $T\in\mathcal{E}$.
It is well-known that the Dirac operator induces a Lipschitz semi-norm: $\|T\|_{1}=\|[D,T]\|$ on $\mathcal{E}$. Hence the analogue of (2.2) in this case is
$d_{E}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1,\|f\|_{1}\leq 1\right\}$ (2.9)
where $\varphi,\psi$ are in $\mathcal{UCP}_{\mathcal{K}}(E)$.
Definition 2.5.
We say that an operator system spectral triple $(\mathcal{E},\mathcal{H},D)$ has the BW property if the metric topology on $\mathcal{UCP}_{\mathcal{K}}\big{(}E\big{)}$ induced by the metric $d_{E}$ coincides with the BW-topology on $\mathcal{UCP}_{\mathcal{K}}\big{(}E\big{)}$.
Lemma 2.6.
Let $(\mathcal{E},\mathcal{H},D)$ be an operator system spectral triple such that $\mathcal{E}$ is finite-dimensional. Then $(\mathcal{E},\mathcal{H},D)$ has the BW property.
Proof.
We are omitting the proof because it is similar to the proof of Lemma 2.3. The compactness argument which involved the Arzelà-Ascoli theorem in the proof of Lemma 2.3 is automatic in the finite-dimensional case because of compactness of the closed unit ball. ∎
Observation 2.7.
Obviously, in the finite dimensional case, one can omit the condition $\|f\|_{1}\leq 1$ in the definition of the metric. The metric could be defined as $d_{E}^{\prime}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1\right\}$. This too would metrize the BW-topology. So, the two metrics $d_{E}$ and $d_{E}^{\prime}$ are equivalent. In other words, the Dirac operator is not necessary to define the metric. We introduce the Dirac operator at this stage because it is essential in formulating the sufficient conditions which will guarantee convergence in the next section.
3. The Gromov–Hausdorff convergence
Now that we know about the metric space structure of $\mathcal{UCP}_{\mathcal{K}}(E)$, we turn to the Gromov-Hausdorff metric $d_{GH}(M_{1},M_{2})$ between two compact metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$.
Definition 3.1.
$d_{GH}(M_{1},M_{2})$ is the infimum of the set
[…]
$f\text{ and }g\text{ are isometric embeddings of }M_{1}\text{ and }M_{2}\text{ into }M\}$
where $d_{H}$ stands for the Hausdorff distance between closed subsets of a compact metric space.
For proving convergence in the Gromov-Hausdorff metric, we use the distortion.
Definition 3.2.
[1, Definition 7.3.1] Let $M_{1}$ and $M_{2}$ be two sets. A total onto correspondence between $M_{1}$ and $M_{2}$ is a relation $\mathfrak{R}\subseteq M_{1}\times M_{2}$ such that for every $x\in M_{1}$ there exists at least one $y\in M_{2}$ with $(x,y)\in\mathfrak{R}$ and similarly for every $y\in M_{2}$ there exists an $x\in M_{1}$ with $(x,y)\in\mathfrak{R}$.
Definition 3.3.
[1, Definition 7.3.10] Let $\mathfrak{R}$ be a total onto correspondence between metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$. The distortion of $\mathfrak{R}$ is defined by
$\operatorname{dis}\mathfrak{R}=\sup\{|{d_{1}(x,x^{\prime})-d_{2}(y,y^{\prime})|}:(x,y),(x^{\prime},y^{\prime})\in\mathfrak{R}\}.$
We depend on Theorem 7.3.25 of [1] which gives the crucial relation that for any two metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$,
$d_{GH}(M_{1},M_{2})=\frac{1}{2}\inf_{\mathfrak{R}}(\operatorname{dis}\mathfrak{R}),$ (3.1)
where the infimum is taken over all total onto correspondences $\mathfrak{R}$ between $M_{1}$ and $M_{2}$.
Definition 3.4.
Let $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ be a sequence of operator system spectral triples and let $(\mathcal{E},\mathcal{H},D)$ be an operator system spectral triple. A $\mathrm{C}^{1}$-approximate complete order isomorphism is a pair $(R_{n},S_{n})$ of linear maps $R_{n}:\mathcal{E}\to\mathcal{E}_{n}$ and $S_{n}:\mathcal{E}_{n}\to\mathcal{E}$ for each $n$, satisfying the following conditions:
  1. (1)
    The maps $R_{n}$ and $S_{n}$ are unital completely positive (ucp) maps.
  2. (2)
[…]
  3. (3)
    There exist sequences $c_{n}$ and $c_{n}^{\prime}$ both converging to zero such that
    $\|S_{n}\circ R_{n}(f)-f\|$ $\leq c_{n}\|f\|_{1},$
    $\|R_{n}\circ S_{n}(h)-h\|$ $\leq c_{n}^{\prime}\|h\|_{1}$
    for all $f\in\mathcal{E}$ and $h\in\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{E}_{n}$.
Since the composition of two ucp maps is again a $\mathrm{ucp}$ map, we can pull back ucp maps as follows:
$R_{n}^{*}:\mathcal{UCP}_{\mathcal{K}}\big{(}E_{n}\big{)}\to\mathcal{UCP}_{\mathcal{K}}(E);\qquad\varphi_{n}\mapsto\varphi_{n}\circ R_{n},$
$S_{n}^{*}:\mathcal{UCP}_{\mathcal{K}}\big{(}E\big{)}\to\mathcal{UCP}_{\mathcal{K}}(E_{n});\qquad\varphi\mapsto\varphi\circ S_{n}.$
The following results transfer [13, Proposition 4] to the setting of $\mathrm{ucp}$ maps.
Proposition 3.5.
If $(R_{n},S_{n})$ is a $C^{1}$-approximate complete order isomorphism for $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$, then
  1. (1)
    For all $\varphi,\psi\in\mathcal{UCP}_{\mathcal{K}}(E)$ we have
    $d_{E_{n}}(\varphi\circ S_{n},\psi\circ S_{n})\leq d_{E}(\varphi,\psi)\leq d_{E_{n}}(\varphi\circ S_{n},\psi\circ S_{n})+2c_{n}.$
  2. (2)
    For all $\varphi_{n},\psi_{n}\in\mathcal{UCP}_{\mathcal{K}}(E_{n})$ we have
    $d_{E}(\varphi_{n}\circ R_{n},\psi_{n}\circ R_{n})\leq d_{E_{n}}(\varphi_{n},\psi_{n})\leq d_{E}(\varphi_{n}\circ R_{n},\psi_{n}\circ R_{n})+2c_{n}^{\prime}.$
Proof.
Let $\varphi,\psi\in\mathcal{UCP}_{\mathcal{K}}(E)$. Note that, for $n\in\mathbb{N}$, $S_{n}:E_{n}\to E$ are Lipschitz contractive as well as contractive; therefore, we have:
$d_{E_{n}}(\varphi\circ S_{n},\psi\circ S_{n})$
$=$ $\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(S_{n}f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E}_{n},\|f\|\leq 1,\|f\|_{1}\leq 1\right\}$
$\leq$ $\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(g)k_{p},k_{p}\rangle\right|:g\in\mathcal{E},\|g\|\leq 1,\|g\|_{1}\leq 1\right\}=d_{E}(\varphi,\psi)$ (3.2)
On the other hand, for all $f\in\mathcal{E}$ with $\|f\|,\,\|f\|_{1}\leq 1$, we have
$\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|$
$\leq\sum_{p=1}^{\infty}\frac{1}{2^{p}}|\varphi(S_{n}(R_{n}(f)))-\psi(S_{n}(R_{n}(f)))|+\sum_{p=1}^{\infty}\frac{1}{2^{p}}|\varphi(f)-\varphi(S_{n}(R_{n}(f)))|$
$+\sum_{p=1}^{\infty}\frac{1}{2^{p}}|\psi(f)-\psi(S_{n}(R_{n}(f)))|$
$\leq d_{E}(\varphi\circ S_{n},\psi\circ S_{n})+2c_{n},$ (3.3)
as $\|\varphi\|=\|\psi\|=1$ and $\|R_{n}(f)\|_{1}\leq\|f\|_{1}\leq 1$. Therefore, (3) and (3) together yield (1). By similar lines of argument, we also prove (2). ∎
Now we state and prove our main result below.
Theorem 1.
Let $n\in\mathbb{N}$. Let $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$, and $(\mathcal{E},\mathcal{H},D)$ be operator system spectral triples having BW-property. If $(R_{n},S_{n})$ is a $C^{1}$-approximate complete order isomorphism for $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$, then $\mathcal{UCP}_{\mathcal{K}}(E_{n})$ converges to $\mathcal{UCP}_{\mathcal{K}}(E)$ in the Gromov–Hausdorff metric.
Proof.
For $n\in\mathbb{N}$, using $R_{n}$, $S_{n}$, $R_{n}^{*}$, and $S_{n}^{*}$, let us define the correspondence $\mathfrak{R}_{n}\subset\mathcal{UCP}_{\mathcal{K}}(E_{n})\times\mathcal{UCP}_{\mathcal{K}}(E)$ by
$\mathfrak{R}_{n}=\left\{(\varphi_{n},R_{n}^{*}(\varphi_{n})):\varphi_{n}\in\mathcal{UCP}_{\mathcal{K}}(E_{n})\right\}\cap\left\{(S_{n}^{*}(\varphi),\varphi):\varphi\in\mathcal{UCP}_{\mathcal{K}}(E)\right\}$
with distortion
$\textup{dis}(\mathfrak{R}_{n})=\sup\left\{\left|d_{E_{n}}(\varphi_{n},\varphi_{n}^{\prime})-d_{E}(\varphi,\varphi^{\prime})\right|:(\varphi_{n},\varphi),(\varphi_{n}^{\prime},\varphi^{\prime})\in\mathfrak{R}_{n}\right\}.$
By (3.1), we have
$d_{GH}(\mathcal{UCP}_{\mathcal{K}}(E_{n}),\mathcal{UCP}_{\mathcal{K}}(E))\leq\frac{1}{2}\operatorname{dis}\mathfrak{R}_{n}.$ (3.4)
It follows from Proposition 3.5 that $\operatorname{dis}\mathfrak{R}_{n}$ is bounded by $2(c_{n}+c_{n}^{\prime})$ and hence converges to zero as $c_{n}$ and $c_{n}^{\prime}$ converge to zero. Therefore
$\lim_{n\to\infty}\,d_{GH}(\mathcal{UCP}_{\mathcal{K}}(E_{n}),\mathcal{UCP}_{\mathcal{K}}(E))=0.$
This complete the proof. ∎
4. Applications
4.1. The Fejér–Riesz operator systems
Let us consider the following spectral triple
$\left(\mathcal{E}:=C^{\infty}(S^{1}),\mathcal{H}:=L^{2}(S^{1}),D:=-i\frac{d}{d\theta}\right).$ (4.1)
Consider the Fejer-Riesz operator system $C(S^{1})_{(n)}$ (cf. Section 1). Since this is an operator subsystem of $E:=C(S^{1})$ it is natural to consider the following spectral triple:
$\left(\mathcal{E}_{n}:=C(S^{1})_{(n)},\mathcal{H}_{n}:=L^{2}(S^{1}),D_{n}:=-i\frac{d}{d\theta}\right).$ (4.2)
We will seek a $C^{1}$-approximate complete order isomorphism $(R_{n},S_{n})$ for the operator system spectral triples $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$, allowing us to apply Theorem 1 to conclude Gromov–Hausdorff convergence of the corresponding spaces of ucp maps.
Let
$R_{n}:C(S^{1})$ $\to C(S^{1})_{(n)}$
$f$ $\mapsto F_{n}\ast f$
where we recall that $F_{n}=\sum_{|k|\leq n-1}(1-|k|/n)e^{ik\theta}$ is the Fejér kernel (cf. [12, Chapter 2]). Also, define
$S_{n}:C(S^{1})_{(n)}$ $\to C(S^{1})$
$f$ $\mapsto f.$
It follows from Lemma 14 and Lemma 15 of [13] that there exist sequences $\{c_{n}\}$ and $\{c_{n}^{\prime}\}$, both converging to 00, such that
$\|S_{n}\circ R_{n}(f)-f\|$ $\leq c_{n}\|f\|_{1},$
$\|R_{n}\circ S_{n}(h)-h\|$ $\leq c_{n}^{\prime}\|h\|_{1}.$
From the definition and the properties of the kernel $F_{n}$, it follows that $R_{n}$ and $S_{n}$ are both unital, positive, contractive, and Lipschitz contractive. Since $C(S^{1})$ is a commutative $C^{*}$-algebra, by [9, Theorems 3.9 and 3.11], $R_{n}$ and $S_{n}$ are ucp maps. Therefore, $(R_{n},S_{n})$ form a $C^{1}$-approximate complete order isomorphism for $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$. Hence, we can apply Theorem 1 to these operator systems.
Corollary 4.1.
$\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(S^{1}))$ in the Gromov-Hausdorff distance.
4.2. The Toeplitz operator systems
Consider the spectral triple $(\mathrm{C}^{\infty}(S^{1}),\mathrm{L}^{2}(S^{1}),-i\frac{d}{d\theta})$ and the truncated spectral triples
$(\mathcal{T}_{n}=\{P_{n}M_{f}|_{\mathcal{H}_{n}}:f\in\mathrm{C}^{\infty}(S^{1})\},\mathcal{H}_{n},-iP_{n}\frac{d}{d\theta}|_{\mathcal{H}_{n}}).$
In the above, $\mathcal{H}_{n}$ is the $n$-dimensional space spanned by $\{e_{1},e_{2},\ldots,e_{n}\}$ where $e_{k}$ is the function $z^{k}$ on the unit circle and $P_{n}$ is the orthogonal projection from $\mathrm{L}^{2}(S^{1})$ onto $\mathcal{H}_{n}$. Note that if $f$ has the Fourier series expansion
$f(z)=\sum_{k=-\infty}^{\infty}a_{k}z^{k},$
then the matrix of $P_{n}M_{f}|_{\mathcal{H}_{n}}$ in the orthonormal basis $\{e_{1},e_{2},\ldots,e_{n}\}$ is the Toeplitz matrix $((a_{k-j}))_{k,j=0}^{n}$. Define $R_{n}$ by sending $f$ in $\mathrm{C}^{\infty}(S^{1})$ to $P_{n}M_{f}|_{\mathcal{H}_{n}}$. The unit circle has a natural action on $\mathcal{T}_{n}$, viz.,
$\alpha_{\theta}\left(((a_{k-j}))_{k,j=0}^{n}\right)=\left(((e^{i(k-j)\theta}a_{i-j}))_{i,j=0}^{n}\right).$
Let $\psi$ be the unit vector $(e_{1}+e_{2}+\cdots+e_{n})/\sqrt{n}$. Define $S_{n}:\mathcal{T}_{n}\rightarrow\mathrm{C}(S^{1})$ by
$(S_{n}(T))(\theta)=\langle\psi,\alpha_{\theta}(T)\psi\rangle.$
See [13] where it was verified that these are $C^{1}$ order isomorphisms. Hence we have the following.
Corollary 4.2.
$\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(S^{1}))$ in the Gromov-Hausdorff distance.
4.3. The $d$-dimensional torus
In this section, $\Omega$ stands for the flat torus $S^{1}\times S^{1}\times\cdots\times S^{1}$ ($d$-times). We consider the operator system spectral triple:
$\left(\mathcal{E}:=\mathrm{C}^{\infty}(\Omega),\mathcal{H}:=\mathrm{L}^{2}(\textit{spb}(\Omega)),D\right).$ (4.3)
Here, $\mathcal{E}$ represents the *-algebra of smooth functions on the torus $\Omega$. These functions act by multiplication on a Hilbert space, denoted by $\mathcal{H}$, which consists of $\mathrm{L}^{2}$-sections of the spinor bundle $\textit{spb}(\Omega)$. We identify $\textit{spb}(\Omega)$ with the trivial bundle $\Omega\otimes V$, where $V:=\mathbb{C}^{\lfloor\frac{d}{2}\rfloor}$, and we write $\mathcal{H}:=\mathrm{L}^{2}(\Omega)\otimes V\cong\mathrm{L}^{2}(\textit{spb}(\Omega))$ for the Hilbert space. The Dirac operator $D$ given by
$D=-i\sum_{\mu=1}^{\lfloor\frac{d}{2}\rfloor}\partial_{\mu}\otimes\gamma^{\mu}$
acts on the dense subspace of smooth sections of the spinor bundle within $\mathcal{H}$.
For an integer $N\geq 0$, let $P_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ of the eigenvalues $\lambda$ with $|\lambda|\leq N$. More concretely, we have $P_{N}\mathcal{H}=\mathrm{span}\{e_{n}\,:\,n\in\mathbb{Z}^{d},\|n\|\leq N\}\otimes V$, with $e_{n}(x):=e^{in\cdot x}$, for all $x\in\Omega$.
Let $\mathcal{T}^{d}_{N}$ be the operator system of $d$-dimensional Toeplitz matrices $T=(t_{k-l})_{k,l\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}}$. Here, $\overline{\mathrm{B}}_{N}^{\mathbb{Z}}$ is the set of points in the integer lattice $\mathbb{Z}^{d}$ that lie inside a ball of radius $N$. The term $t_{k-l}$ represents the inner product $\langle e_{k},Te_{l}\rangle$. Specifically, for $d=1$, the operator system $\mathcal{T}^{1}_{N}$ corresponds to the set of Toeplitz matrices of size $(2N+1)\times(2N+1)$, which were studied in detail in [2] and [6].
The spectral projection $P_{N}$ gives rise to the following operator system spectral triple:
$\left(\mathcal{E}_{N}:=\mathcal{T}^{d}_{N},\mathcal{H}_{N}:=P_{N}\mathcal{H},D_{N}:=P_{N}D|_{\mathcal{H}_{N}}\right),$ (4.4)
where $\mathcal{H}$ and $D$ are given by (4.3). Note that the truncations for the torus are different from the truncation for the unit circle discussed above. We recall the following from [8].
Define
$R_{N}:$ $~{}\mathcal{E}\to\mathcal{E}_{N}$
$f\mapsto P_{N}fP_{N},$
$S_{N}:$ $~{}\mathcal{E}_{N}\to\mathcal{E}$
$T\mapsto\frac{1}{\mathcal{N}_{\mathrm{B}}(N)}\mathrm{Tr}\left(|\psi\rangle\langle\psi|\alpha(T)\right).$
Here, $\alpha$ is the $\Omega$-action on $\mathcal{E}_{N}$ given by
$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}},$ (4.5)
where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$, the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}}e_{n}$, and $\mathcal{N}_{\mathrm{B}}(N):=\#\overline{\mathrm{B}}_{N}\cap\mathbb{Z}^{d}$ is the number of $\mathbb{Z}^{d}$-lattice points in the closed ball of radius $N$. By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have BW-property.
It is shown in the proof of [8, Theorem 3.10] that the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$. By [9, Theorems 3.9 and 3.11], $R_{N}$ and $S_{N}$ are ucp maps, and hence $(R_{N},S_{N})$ form a $C^{1}$-approximate complete order isomorphism. Therefore, applying Theorem 1, we have the following convergnce result.
Corollary 4.3.
Let $d\in\mathbb{N}$. Then $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{N}^{d})$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$ in the Gromov-Hausdorff distance.

### v2

@@
 Keywords: Operator system, UCP maps, Spectral truncation, Gromov–Hausdorff convergence
-Acknowledgement: The authors are thankful to Walter van Suijlekom whose India visit under the Jubilee Chair Professorship of the Indian of Academy of Sciences was pivotal for this paper.
 1. Introduction
@@
 More generally, in [16], general conditions were found on operator system spectral triples that guarantee the Gromov-Hausdorff convergence of state spaces.
-Given an operator system $E$ and a (complex, separable) Hilbert space $\mathcal{K}$, let $\mathcal{UCP}_{\mathcal{K}}(E)$ denote the collection of all unital completely positive (UCP) maps from $E$ into the algebra of all bounded operators $\mathcal{B}(\mathcal{K})$.
+Given a closed operator system $E$ and a (complex, separable) Hilbert space $\mathcal{K}$, let $\mathcal{UCP}_{\mathcal{K}}(E)$ denote the collection of all unital completely positive (UCP) maps from $E$ into the algebra of all bounded operators $\mathcal{B}(\mathcal{K})$.
 We shall show in this note that the general conditions due to van Suijlekom mentioned above actually guarantee much more, viz., the Gromov-Hausdorff convergence of sequences of sets of UCP maps.
@@
 The main result is Theorem 1 where we consider a sequence of operator system spectral triples $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ related to an operator system spectral triple $(\mathcal{E},\mathcal{H},D)$ by a $\mathrm{C}^{1}$-approximate complete order isomorphism.
-Section 4 contains applications of the theorem to examples like the ones mentioned above.
+Section 4 contains applications of the theorem to examples like the ones mentioned above as well as for polyhedral truncations motivated by works of Travaglini [17].
 The $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ are often called spectral truncations of $(\mathcal{E},\mathcal{H},D)$.
@@
 For fixed vectors $x$ in $X$ and $y$ in $Y$, define a linear functional $x\otimes y$ on $\mathcal{B}(X,Y^{*})$ by $x\otimes y(L)$ = $L(x)(y)$.
-Let $Z$ denote the closed linear span in $\mathcal{B}(X,Y^{*})$ of these elementary tensors.
+Let $Z$ denote the closed linear span in $(\mathcal{B}(X,Y^{*}))^{*}$ of these elementary tensors.
 Then $\mathcal{B}(X,Y^{*})$ is isometrically isomorphic to $Z^{*}$ with the duality given by
@@
 The weak topology that is induced on $\mathcal{B}(X,Y^{*})$ by this identification is called the BW topology ((((for bounded weak)))).
-If $Y$ is the algebra $\mathcal{B}(\mathcal{H})$ of bounded operators on a Hilbert space $\mathcal{H}$, then a bounded net $L_{\lambda}$ in $\mathcal{B}(X,\mathcal{B}(H))$ converges in the BW topology to $L$ if and only if $\langle L_{\lambda}(x)h,h^{\prime}\rangle$ converges to $\langle L(x)h,h^{\prime}\rangle$ for all $h$, $h^{\prime}$ in $\mathcal{H}$ and $x$ in $X$.
+If $Y^{*}$ is the algebra $\mathcal{B}(\mathcal{H})$ of bounded operators on a Hilbert space $\mathcal{H}$, then a bounded net $L_{\lambda}$ in $\mathcal{B}(X,\mathcal{B}(H))$ converges in the BW topology to $L$ if and only if $\langle L_{\lambda}(x)h,h^{\prime}\rangle$ converges to $\langle L(x)h,h^{\prime}\rangle$ for all $h$, $h^{\prime}$ in $\mathcal{H}$ and $x$ in $X$.
 Naturally, $\mathcal{UCP}_{\mathcal{K}}(E)$ inherits the BW-topology from $\mathcal{B}(E,\mathcal{B}(K))$.
-It is known that $\mathcal{UCP}_{\mathcal{K}}(E)$ is compact in the BW-topology [9, Theorem 7.4].
+It is known that $\mathcal{UCP}_{\mathcal{K}}(E)$ is compact in the BW-topology [12, Theorem 7.4].
 2.2. A metric for the BW-topology when $E=C(\Omega)$.
-Let $(\Omega,d_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega})$ a compact metric space. Consider the set of all Lipschitz functions
-$\text{Lip}(\Omega):=\{f:\Omega\to\mathbb{C}\text{ such that }\|f\|_{1}:=\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{d_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega}(x,y)}<\infty\}.$ (2.1)
+Let $(\Omega,d_{\Omega})$ be a compact metric space. Consider the set of all Lipschitz functions
+$\text{Lip}(\Omega):=\{f:\Omega\to\mathbb{C}\text{ such that }\|f\|_{1}:=\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{d_{\Omega}(x,y)}<\infty\}.$ (2.1)
 Let $\{k_{n}\in\mathcal{K}:n\in\mathbb{N}\}$ be a fixed dense subset of the closed unit ball of $\mathcal{K}$.
@@
 Proof.
-Clearly, symmetry and that the triangle inequality hold.
+Clearly, symmetry and the triangle inequality hold.
 Also, the implication $\varphi=\psi\implies d(\varphi,\psi)=0$ is immediate.
-Suppose now $d(\varphi,\psi)=0$.
-Then for all $f\in\text{Lip}(\Omega)$ and for all $p$,
-$\left\langle(\varphi-\psi)(f)k_{p},k_{p}\right\rangle=0\implies\left\langle(\varphi-\psi)(f)k,k\right\rangle=0$ (2.3)
-for all $k\in\mathcal{K}$. For $h,k\in\mathcal{K}$, (2.3) and the polarization identity
+Suppose now that $d(\varphi,\psi)=0$.
+Then for all $f\in\text{Lip}(\Omega)$ and for all $p$, we have $\left\langle(\varphi-\psi)(f)k_{p},k_{p}\right\rangle=0$.
+By denseness, this implies that
+$\left\langle(\varphi-\psi)(f)k,k\right\rangle=0\text{ for all }k\in\mathcal{K}.$ (2.3)
+For $h,k\in\mathcal{K}$, (2.3) and the polarization identity
 $\left\langle(\varphi-\psi)(f)h,k\right\rangle=\sum_{l=0}^{3}\frac{i^{l}}{4}\left\langle(\varphi-\psi)(f)(h+i^{l}k),h+i^{l}k\right\rangle$
-together yield $\left\langle(\varphi-\psi)(f)h,k\right\rangle=0$. Hence $(\varphi-\psi)(f)=0$ for all $f\in\text{Lip}(\Omega)$, and consequently $\varphi=\psi$. This completes the proof.
+together yield $\left\langle(\varphi-\psi)(f)h,k\right\rangle=0$. Hence $(\varphi-\psi)(f)=0$ for all $f\in\text{Lip}(\Omega)$, and consequently $\varphi=\psi$. This completes the proof. ∎
+The following is probably known. We give a proof because we could not find a suitable reference in the literature.
 Lemma 2.3.
@@
 Then there exist $\varepsilon>0$ and an infinite subset $\hat{\Lambda}\subseteq\Lambda$ such that $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for all $\lambda\in\hat{\Lambda}$.
-Without loss of generality, we may assume that $\hat{\Lambda}=\mathbb{N}$.
+Invoking the Axiom of Choice, we may assume that $\hat{\Lambda}=\mathbb{N}$.
 Thus, we have $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for $\lambda\in\mathbb{N}$.
@@
 $\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|>2\varepsilon.$ (2.4)
-Since $\|f_{\lambda}\|_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}\leq 1$, so $\{f_{\lambda}\}$ is a sequence of equicontinuous functions.
-Next, we also note that the series in (2.4) is convergent uniformly in $\lambda$ as
-$\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|\leq 2.$
-Hence, for the given $\varepsilon>0$, there exist a $N_{0}\in\mathbb{N}$ such that
-$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p-1}}<\varepsilon.$
-Consequently, for every $\lambda$
-$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|<\varepsilon.$ (2.5)
-Therefore, (2.4) and (2.5) together give
-$\sum_{p=1}^{N_{0}}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|>\varepsilon.$ (2.6)
-Since there are only finitely many $p$ in (2.6), we can choose a $p_{0}\in\{1,\ldots,N_{0}\}$ such that
+Since $\|f_{\lambda}\|_{1}\leq 1$, so $\{f_{\lambda}\}$ is a sequence of equicontinuous functions. Next, in view of
+$\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|\leq 2$
+and the fact that for a given $\varepsilon>0$, there exist an $N_{0}\in\mathbb{N}$ such that
+$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p-1}}<\varepsilon,$
+we have
+$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|<\varepsilon$ (2.5)
+for every $\lambda$. Therefore, (2.4) and (2.5) together give
+$\sum_{p=1}^{N_{0}}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|>\varepsilon\;\;\text{ for every $\lambda$}.$ (2.6)
+Since there are only finitely many $p$ in (2.6), we can choose a $p_{0}\in\{1,\ldots,N_{0}\}$ such that after passing to a subsequence and continuing to denote the subsequence by $\lambda$, we have
 $\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p_{0}},k_{p_{0}}\right\rangle\right|>\frac{\varepsilon}{N_{0}}.$ (2.7)
@@
 It is well-known that the Dirac operator induces a Lipschitz semi-norm: $\|T\|_{1}=\|[D,T]\|$ on $\mathcal{E}$. Hence the analogue of (2.2) in this case is
-$d_{E}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1,\|f\|_{1}\leq 1\right\}$ (2.9)
+$d_{E}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1,\|f\|_{1}\leq 1\right\}$
 where $\varphi,\psi$ are in $\mathcal{UCP}_{\mathcal{K}}(E)$.
@@
 Observation 2.7.
-Obviously, in the finite dimensional case, one can omit the condition $\|f\|_{1}\leq 1$ in the definition of the metric.
-The metric could be defined as $d_{E}^{\prime}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1\right\}$.
+Obviously, in the finite dimensional case, one can omit the condition $\|f\|_{1}\leq 1$ in the definition of the metric. The metric could be defined as
+$d_{E}^{\prime}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1\right\}.$
 This too would metrize the BW-topology.
@@
 3. The Gromov–Hausdorff convergence
-Now that we know about the metric space structure of $\mathcal{UCP}_{\mathcal{K}}(E)$, we turn to the Gromov-Hausdorff metric $d_{GH}(M_{1},M_{2})$ between two compact metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$.
+In this section, we shall consider the usual Gromov-Hausdorff metric $d_{GH}(M_{1},M_{2})$ between two compact metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$ and apply it to measure distances between operator system spectral triples.
+For various notions of distances between Lip-normed operator systems and inter-relations between them, we refer the reader to section 3 of [9].
 Definition 3.1.
@@
     $\|R_{n}\circ S_{n}(h)-h\|$ $\leq c_{n}^{\prime}\|h\|_{1}$
-    for all $f\in\mathcal{E}$ and $h\in\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{E}_{n}$.
+    for all $f\in\mathcal{E}$ and $h\in\mathcal{E}_{n}$.
 Since the composition of two ucp maps is again a $\mathrm{ucp}$ map, we can pull back ucp maps as follows:
@@
 $S_{n}^{*}:\mathcal{UCP}_{\mathcal{K}}\big{(}E\big{)}\to\mathcal{UCP}_{\mathcal{K}}(E_{n});\qquad\varphi\mapsto\varphi\circ S_{n}.$
-The following results transfer [13, Proposition 4] to the setting of $\mathrm{ucp}$ maps.
+The following results transfer [16, Proposition 4] to the setting of $\mathrm{ucp}$ maps.
 Proposition 3.5.
@@
 By (3.1), we have
-$d_{GH}(\mathcal{UCP}_{\mathcal{K}}(E_{n}),\mathcal{UCP}_{\mathcal{K}}(E))\leq\frac{1}{2}\operatorname{dis}\mathfrak{R}_{n}.$ (3.4)
+$d_{GH}(\mathcal{UCP}_{\mathcal{K}}(E_{n}),\mathcal{UCP}_{\mathcal{K}}(E))\leq\frac{1}{2}\operatorname{dis}\mathfrak{R}_{n}.$
 It follows from Proposition 3.5 that $\operatorname{dis}\mathfrak{R}_{n}$ is bounded by $2(c_{n}+c_{n}^{\prime})$ and hence converges to zero as $c_{n}$ and $c_{n}^{\prime}$ converge to zero.
@@
 This complete the proof. ∎
+A different proof for finite-dimensional $\mathcal{K}$ can be given using [10, Proposition 5.19] of Leimbach.
+This was communicated to us by him.
+He showed a stronger result, viz., an estimate on the complete Gromov-Hausdorff distance ([8, Definition 3.2]).
+The idea of his proof stems from Proposition 2.14 of [7] which uses the notion of bridge introduced by Rieffel [13].
 4. Applications
@@
 Let us consider the following spectral triple
-$\left(\mathcal{E}:=C^{\infty}(S^{1}),\mathcal{H}:=L^{2}(S^{1}),D:=-i\frac{d}{d\theta}\right).$ (4.1)
+$\left(\mathcal{E}:=C^{\infty}(S^{1}),\mathcal{H}:=L^{2}(S^{1}),D:=-i\frac{d}{d\theta}\right).$
 Consider the Fejer-Riesz operator system $C(S^{1})_{(n)}$ (cf. Section 1). Since this is an operator subsystem of $E:=C(S^{1})$ it is natural to consider the following spectral triple:
-$\left(\mathcal{E}_{n}:=C(S^{1})_{(n)},\mathcal{H}_{n}:=L^{2}(S^{1}),D_{n}:=-i\frac{d}{d\theta}\right).$ (4.2)
+$\left(\mathcal{E}_{n}:=C(S^{1})_{(n)},\mathcal{H}_{n}:=L^{2}(S^{1}),D_{n}:=-i\frac{d}{d\theta}\right).$
+That these two spectral triples have the BW property is a consequence of results of section 2.
 We will seek a $C^{1}$-approximate complete order isomorphism $(R_{n},S_{n})$ for the operator system spectral triples $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$, allowing us to apply Theorem 1 to conclude Gromov–Hausdorff convergence of the corresponding spaces of ucp maps.
@@
 $f$ $\mapsto F_{n}\ast f$
-where we recall that $F_{n}=\sum_{|k|\leq n-1}(1-|k|/n)e^{ik\theta}$ is the Fejér kernel (cf. [12, Chapter 2]). Also, define
+where we recall that $F_{n}=\sum_{|k|\leq n-1}(1-|k|/n)e^{ik\theta}$ is the Fejér kernel (cf. [15, Chapter 2]). Also, define
 $S_{n}:C(S^{1})_{(n)}$ $\to C(S^{1})$
@@
 From the definition and the properties of the kernel $F_{n}$, it follows that $R_{n}$ and $S_{n}$ are both unital, positive, contractive, and Lipschitz contractive.
-Since $C(S^{1})$ is a commutative $C^{*}$-algebra, by [9, Theorems 3.9 and 3.11], $R_{n}$ and $S_{n}$ are ucp maps.
+Since $C(S^{1})$ is a commutative $C^{*}$-algebra, by [12, Theorems 3.9 and 3.11], $R_{n}$ and $S_{n}$ are ucp maps.
 Therefore, $(R_{n},S_{n})$ form a $C^{1}$-approximate complete order isomorphism for $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$.
@@
 These functions act by multiplication on a Hilbert space, denoted by $\mathcal{H}$, which consists of $\mathrm{L}^{2}$-sections of the spinor bundle $\textit{spb}(\Omega)$.
-We identify $\textit{spb}(\Omega)$ with the trivial bundle $\Omega\otimes V$, where $V:=\mathbb{C}^{\lfloor\frac{d}{2}\rfloor}$, and we write $\mathcal{H}:=\mathrm{L}^{2}(\Omega)\otimes V\cong\mathrm{L}^{2}(\textit{spb}(\Omega))$ for the Hilbert space.
+We identify $\textit{spb}(\Omega)$ with the trivial bundle $\Omega\otimes V$, where $V:=\mathbb{C}^{d}$, and we write $\mathcal{H}:=\mathrm{L}^{2}(\Omega)\otimes V\cong\mathrm{L}^{2}(\textit{spb}(\Omega))$ for the Hilbert space.
 The Dirac operator $D$ given by
-$D=-i\sum_{\mu=1}^{\lfloor\frac{d}{2}\rfloor}\partial_{\mu}\otimes\gamma^{\mu}$
+$D=-i\sum_{\mu=1}^{d}\partial_{\mu}\otimes\gamma^{\mu}$
 acts on the dense subspace of smooth sections of the spinor bundle within $\mathcal{H}$.
-For an integer $N\geq 0$, let $P_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ of the eigenvalues $\lambda$ with $|\lambda|\leq N$.
+4.3.1. Spherical truncation
+For an integer $N\geq 0$, let $P_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ with $|\lambda|\leq N$.
 More concretely, we have $P_{N}\mathcal{H}=\mathrm{span}\{e_{n}\,:\,n\in\mathbb{Z}^{d},\|n\|\leq N\}\otimes V$, with $e_{n}(x):=e^{in\cdot x}$, for all $x\in\Omega$.
@@
 $\left(\mathcal{E}_{N}:=\mathcal{T}^{d}_{N},\mathcal{H}_{N}:=P_{N}\mathcal{H},D_{N}:=P_{N}D|_{\mathcal{H}_{N}}\right),$ (4.2)
-where $\mathcal{H}$ and $D$ are given by (4.3). Note that the truncations for the torus are different from the truncation for the unit circle discussed above. We recall the following from [8].
+where $\mathcal{H}$ and $D$ are given by (4.1).
+We call the above truncation (4.2) the spherical truncation of the (4.1).
+Note that the truncations for the torus are different from the truncation for the unit circle discussed above.
+We recall the following from [11].
 Define
@@
 Here, $\alpha$ is the $\Omega$-action on $\mathcal{E}_{N}$ given by
-$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}},$ (4.5)
-where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$, the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}}e_{n}$, and $\mathcal{N}_{\mathrm{B}}(N):=\#\overline{\mathrm{B}}_{N}\cap\mathbb{Z}^{d}$ is the number of $\mathbb{Z}^{d}$-lattice points in the closed ball of radius $N$.
-By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have BW-property.
-It is shown in the proof of [8, Theorem 3.10] that the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$.
-By [9, Theorems 3.9 and 3.11], $R_{N}$ and $S_{N}$ are ucp maps, and hence $(R_{N},S_{N})$ form a $C^{1}$-approximate complete order isomorphism.
+$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}},$
+where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$, the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}}e_{n}$, and $\mathcal{N}_{\mathrm{B}}(N):=|\overline{\mathrm{B}}_{N}\cap\mathbb{Z}^{d}|$ is the number of $\mathbb{Z}^{d}$-lattice points in the closed ball of radius $N$.
+By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have the BW-property.
+It is shown in the proof of [11, Theorem 3.10] that the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$.
+By [12, Theorems 3.9 and 3.11], $R_{N}$ and $S_{N}$ are ucp maps, and hence the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate complete order isomorphism.
 Therefore, applying Theorem 1, we have the following convergnce result.
@@
 Let $d\in\mathbb{N}$. Then $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{N}^{d})$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$ in the Gromov-Hausdorff distance.
+4.3.2. Polyhedral truncation
+This section is inspired by [17].
+Definition 4.4.
+Let $r,d\in\mathbb{N}$.
+A polyhedron $\widetilde{\Delta}(r)$ in $\mathbb{R}^{d}$ is the convex hull of a finite set $W=\{\bm{a}_{1},\ldots,\bm{a}_{r}\}\subset\mathbb{Z}^{d}$, where $W$ is minimal, meaning that it coincides with the set of vertices of $\widetilde{\Delta}(r)$.
+The restriction $\Delta(r)=\widetilde{\Delta}(r)\cap\mathbb{Z}^{d}$ is called a summability polyhedron if $\widetilde{\Delta}(r)$ contains the origin strictly in its interior.
+For any $N\in\mathbb{N}$, we denote by $\widetilde{\Delta}_{N}(r)$ the convex hull of the set $\{N\bm{a}_{1},\ldots,N\bm{a}_{r}\}$ and define $\Delta_{N}(r)=\widetilde{\Delta}_{N}(r)\cap\mathbb{Z}^{d}$.
+When the polyhedron is fixed, that is, $r$ and $W$ are fixed, we simply write $\Delta$ and $\Delta_{N}$ instead of $\Delta(r)$ and $\Delta_{N}(r)$, respectively.
+We call $\Delta_{N}$ the $N$-dilation of $\Delta$.
+Let $\Delta$ be a summability polyhedron and let $\Delta_{N}$ be the $N$-dilation of $\Delta$ for an integer $N\geq 1$.
+Let $Q_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ with $\lambda\in\Delta_{N}$, i.e., $Q_{N}\mathcal{H}=\mathrm{span}\{e_{\lambda}\,:\lambda\in\Delta_{N}\}\otimes V$, with $e_{n}(x):=e^{in\cdot x}$, for all $x\in\Omega$.
+Let $\mathcal{T}^{d}_{N}(PH)$ (PH stands for polyhedron) be the operator system of $d$-dimensional polyhedral Toeplitz matrices $T=(t_{k-l})_{k,l\in\Delta_{N}}$.
+This is a $|\Delta_{N}|\times|\Delta_{N}|$ matrix with entries indexed by the elements of $\Delta_{N}$.
+The term $t_{k-l}$ represents the inner product $\langle e_{k},Te_{l}\rangle$.
+The spectral projection $Q_{N}$ defines the operator system spectral triple:
+$\left(\mathcal{E}_{N}:=\mathcal{T}^{d}_{N}(PH),\mathcal{H}_{N}:=Q_{N}\mathcal{H},D_{N}:=Q_{N}D|_{\mathcal{H}_{N}}\right),$ (4.3)
+where $\mathcal{H}$ and $D$ are given by (4.1). We call the above truncation (4.3) a polyhedral truncation of (4.1). Define
+$\rho_{N}:$ $~{}\mathcal{E}\to\mathcal{E}_{N}$
+$f\mapsto Q_{N}fQ_{N},$
+$\sigma_{N}:$ $~{}\mathcal{E}_{N}\to\mathcal{E}$
+$T\mapsto\frac{1}{|\Delta_{N}|}\mathrm{Tr}\left(|\psi\rangle\langle\psi|\alpha(T)\right).$
+Here, $\alpha$ is the $\Omega$-action on $\mathcal{E}_{N}$ given by
+$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\Delta_{N}},$
+where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ is such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$ and the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\Delta_{N}}e_{n}$.
+By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have the BW-property.
+Using the approach of [11], we show below that the pair $(\rho_{N},\sigma_{N})$ forms a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$.
+Lemma 4.5.
+The maps $\rho_{N}$ and $\sigma_{N}$ are unital, completely positive and $\mathrm{C}^{1}$-contractive.
+Proof.
+The proof is similar to [11, Lemma 3.1 and Lemma 3.2], along with an application of [12, Theorems 3.9 and 3.11] as $C^{\infty}(\Omega)$ is commutative. ∎
+Proposition 4.6.
+The pair $(\rho_{N},\sigma_{N})$ forms a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ given by (4.3) and (4.1), respectively.
+Proof.
+By a computation similar to [11, Lemma 3.3], we have
+$\sigma_{N}\circ\rho_{N}(f)(x)=\sum_{n\in\Delta_{N}-\Delta_{N}}\frac{|\Delta_{N}\cap(\Delta_{N}+n)|}{|\Delta_{N}|}\widehat{f}(n)e_{n}(x)$
+$\rho_{N}\circ\sigma_{N}(T)=\left(\frac{|\Delta_{N}\cap(\Delta_{N}+m-n)|}{|\Delta_{N}|}t_{m-n}\right)_{m,n\in\Delta_{N}}$
+For each $N\in\mathbb{N}$, define $K_{2N}$ on $\Omega$ by
+$K_{2N}(x)=$ $\sum_{n,m\in\Delta_{N}}\frac{1}{|\Delta_{N}|}e_{n-m}(x)$
+$=$ $\sum_{n\in\Delta_{N}-\Delta_{N}}\frac{|\Delta_{N}\cap(\Delta_{N}+n)|}{|\Delta_{N}|}e_{n}(x).$
+Then it is easy to check that
+$\sigma_{N}\circ\rho_{N}(f)(x)=(K_{2N}*f)(x).$
+A careful examination of the proof of [11, Theorem 3.10] suggests that, to demonstrate that $(\rho_{N},\sigma_{N})$ form a $C^{1}$-approximate order isomorphism, it suffices to show that the family $\{K_{2N}\}_{N\in\mathbb{N}}$ serves as a good kernel.
+In [17], $K_{2N}$ is called the polyhedron Fejér kernel.
+From [17, Theorems 2, 3], it follows that:
+  1. (1)
+    $K_{2N}\geq 0$,
+  2. (2)
+[…]
+  3. (3)
+    $\hat{K}_{2N}\to 1$ pointwise on $\mathbb{Z}^{d}$.
+Next, we need that the mass of $K_{2N}$ is concentrated near the origin.
+But this can be proved by mimicking the proof of [11, Lemma 3.8] and we omit the details.
+Hence $\{K_{2N}\}$ is good kernel.
+This concludes the proof.
+∎
+Corollary 4.7.
+Let $d\in\mathbb{N}$. Then $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{N}^{d}(PH))$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$ in the Gromov-Hausdorff distance.
+Remark 4.8.
+As one more example of Gromov-Hausdorff convergence of CP maps, consider $V_{n}$, the $n$-dimensional irreducible representation of $SU(2)$.
+Let $S^{2}$ be the $2$-dimensional unit sphere.
+Then $\mathcal{UCP}_{\mathcal{K}}(L(V_{n}))$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(S^{2}))$.
+This is a consequence of non-trivial results of van Suijlekom in section 3.3 of [16] and our Theorem 1.
+van Suijlekom produces the $C^{1}$-approximate order isomorphism and hence Theorem 1 can be applied.
+Acknowledgement: The authors are thankful to Walter van Suijlekom, whose India visit under the Jubilee Chair Professorship of the Indian of Academy of Sciences was pivotal for this paper, as well as M.
+Leimbach for his comments.
+The first author is supported by a J C Bose Fellowship JCB/2021/000041 of SERB.
+The second author is supported by the Prime Minister’s Research Fellowship.
+The third author is supported by NBHM post-doctoral fellowship.
+Declarations
+On behalf of all authors, the corresponding author states that there is no conflict of interest.
+This manuscript has no associated data.